EDITORIAL

Terry Goodman

With the end of the semester approaching, this is always a busy time for us. You will note in this issue that we have dedicated more pages to articles. We are trying to process more efficiently a backlog of manuscripts that we have accepted for publication. Since we are including more articles in this issue we will not have a Problems section. This section will return in our fall issue.

I just spent some time browsing the Web (I planned to spend 30 minutes and ended up spending over an hour). I find that perusing the Web sites for various professional organizations helps me keep track of the information they have available. I have summarized below information currently found on these sites that I believe will be of interest to many of our readers.

http://www.maa.org The Mathematical Association of America’s home page. Here you can find information about this summer’s MathFest, two new MAA books, Euler: The Master of Us All and A Course in Mathematical Modeling, and a proposed new mission statement.

http://www.ams.org The American Mathematics Society’s home page. This site contains information about the new AMS Global Directory of Preprint and e– Print Servers, the upcoming Erdos Memorial Lecture in Denton, Texas, a listing of resources for high school and college students, and summaries of articles on recent developments in mathematics.

http://www.nctm.org The National Council of Teachers of Mathematics’ home page. Here you will find a draft of the Standards 2000 document, a report on teaching and learning mathematics in poor communities, a link to the NCATE Elementary Teacher Program Standards, and an online discussion of student–centered learning and teacher preparation.

http://www.nsf.gov The National Science Foundation’s home page. Here, under Mathematics and Physical Sciences, you will find information concerning NSF’s Fastlane proposal implementation plan, information concerning the agency’s various programs, and a selection of news articles related to this program area.

http://www.learningfirst.org The Learning First Alliance’s home page. Here you will find a copy of Every Child Mathematically Proficient, an action plan of the Alliance, an organization of 12 leading education associations.